OpenStudy (anonymous):

9x-17=6x-8, could someone help me solve and tell me when do i need to change the problem from addition to subtraction

OpenStudy (anonymous):

You need to subtract 6x from both sides first

OpenStudy (anonymous):

ok now 3x-17=-8...add 17 to both sides?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You always do the opposite sign of what you are trying to get rid of when you add or subtract. Hint.. When x's are on both sides I ALWAYS get rid of the smallest x, that way you always have a POSITIVE x to work with after that.

OpenStudy (anonymous):

x=3?

OpenStudy (anonymous):

o btw thank u very much for ur help

OpenStudy (anonymous):

You are correct !!!!

OpenStudy (anonymous):

ur awesome ur a life saver

OpenStudy (anonymous):

-4x+3=6x-7 ........x=-2? do u mind telling me if thats right

OpenStudy (anonymous):

ur awesome ur a life saver